# A Domain-specific Language and Architecture for Detecting Process Activities from Sensor Streams in IoT

### v1

@@ 5.3 Discussion
The results of the evaluation show that we are able to reach good to high quality activity detection levels for the majority of activities by incorporating domain expertise regarding sensor change patterns and conditions into the detection apps. The knowledge of domain experts using the DSL hereby plays an essential role to influence and improve the quality of detections by anticipating potential variations in the sensor data. For some types of activities, the results also show a drop in the detection quality, which is usually due to sensor data not being sufficiently available to distinguish similar activities from each other or to detect them at all. Sections 5.3.1 and 5.3.2 discuss these aspects for the two investigated scenarios in more detail. In general, we face a trade-off between increasing the privacy-invasiveness of the monitoring setup using more capable sensors (e.g., cameras) to increase the quality of activity detections, but then also requiring more complex pre-processing steps (e.g., using computer vision) for abstracting the sensor data to be used in the pattern specifications.
When writing the Radiant applications for the evaluations, the domain experts highlighted the integrated support through the IDE for efficient “programming” of the activity detection patterns. Compared to ML-based models and inference, the proposed pattern-based approach is very light-weight and facilitates the modification and understanding of the sensor patterns leading to the process-level events. Nevertheless, ML-based sensor processing components can easily be integrated into the Radiant applications assuming that they emit abstracted sensors events to a message broker according to the proposed runtime architecture (cf. Section 4.6). In contrast to related approaches only working in offline analysis settings, the underlying CEP engine is specifically designed to process streams of sensor events at runtime, allowing us to provide online feedback regarding the detected activities. Note that we do not expect the activity detections to be used for real-time control of CPS or their components, but to analyze process executions and provide feedback in near real-time to users (e.g., a warning in case the healthcare worker forgot to sanitize the hands before performing a specific treatment activity, indicating non-conformance with the process model [18]) or to other information systems for further analysis or storage.
@@
The seamless workflow from writing a Radiant application to the live system detecting activities from IoT sensor data allows us to conduct efficient evaluations of Radiant applications. Given recorded IoT data and a ground truth log, we are able to quickly inform the domain expert about the detection qualities of the composed applications, potential ambiguities and the need for improvements, e.g., by integrating additional patterns or installing additional sensors to increase the monitoring quality, if possible. Furthermore, in contrast to many black-box ML models, the pattern-based approach makes detection logic explicit and understandable by humans. The availability of a pre-existing, recorded IoT sensor log and associated ground truth of activity executions could also help the domain expert with identifying relevant patterns in the sensor data that should be part of the Radiant applications in a reverse engineering approach as described in [14]. Once the domain expert decides that the quality of the activity detection based on a corresponding Radiant application is sufficient for the specific use case, a Siddhi app can be generated and deployed to the runtime system which enables the sensor-based monitoring of the activity and eliminates the need for additional heavy-weight or manual process monitoring solutions (e.g.,a BPM system) [7].
5.3.1 Smart Manufacturing
@@
The results of the activity detections in the smart manufacturing scenario (cf. Figure 9) show that we can achieve high levels of correct detections when there is only one specific type of activity executed on a production station (e.g., in the sorting machine, oven, and milling machine). Here we are also able to handle variations in sensor patterns quite well via underspecifications and alternative patterns (e.g., in the sorting machine). The vacuum gripper robot (VGR) is a central transport entity, which exhibits similar sensor patterns in its movements for different types of activities. Activities can be mostly distinguished based on the x,y,z-target coordinates of the robot, which need to be discretized by the expert first to represent specific locations. A notable station is the high-bay warehouse (HBW) where there are two different types of activities executed (store and unload). The relevant movement patterns are mostly identical which leads to rather low quality detections. Here we miss a sensor to determine if the buckets being stored or unloaded contain a workpiece or not, which discerns both activities. Moreover, these two activities generally exhibit high amounts of variability as they are executed differently based on the locations of the buckets in the 3x3 storage matrix. Adding more conditions, intermediate and disjunctive patterns might help to increase accuracy here but will also lead to complex Radiant applications and overfitting. Alternatively, the granularity of the activities can be adjusted to, e.g., consider the storage and unloading of an item in each slot of the warehouse as a distinct activity type (e.g., store/unload in/from bucket 1).
@@
Note that in general, the Event Analysis metrics are better for all detections than the Two Set metrics, which also consider the specific points in time when the activity executions happened [70]. These misalignments can be attributed to the software stack controlling the individual production stations [68]. Here we use a layered service-based architecture which introduces some latencies between service executions, logged by the BPM system, and the actual executions in the IoT system that manifest themselves as changes in its sensors and actuators. These latencies might be unintentional (e.g., due to network communication) or intentional (e.g., the services might perform some calculations or data processing before starting to manipulate the sensors and actuators).
5.3.2 Smart Healthcare
@@
The results of the activity detection in the healthcare scenario (cf. Figure 10) show a lower quality of detections. As described in Section 2.2, the activities in this scenario are almost completely manual and strongly affected by variations in the underlying IoT sensor data, which cannot be completely handled by discretizations and alternative detection patterns. While we achieve very good detections for the Sanitize hands activity relying on sensor patterns in the scale and distance sensor in front of it, activities related to the interaction with the donor (e.g., apply/remove tourniquet, insert needle, monitor patient) are harder to detect and distinguish from each other as they are quite similar and rely on the same sensors and similar patterns (cf. simultaneous detection of low-level patterns for activities in Figure 7). For these activities, we currently consider using cameras in combination with the existing sensors for disambiguation. Moreover, the starting and stopping of the blood drawing machine rely on the same button to be pressed, thus we cannot distinguish them from each other purely on the sensor data. Currently, we also investigate considering the process context (e.g., which activities happened before) for further disambiguation. Moreover, similar to the event log serving as ground truth for the smart manufacturing settings, we also observe timeshifts when aligning the start and end of activities for the Two Set metrics in the smart healthcare setting, which leads to a decreased performance compared to Event Analysis. These shifts can be attributed to the creation of the ground truth event log, which is based on manual monitoring and tracking by an external observer of the executed processes. Here, the detection of start and end patterns from the IoT sensor data might not always be aligned with the manual tracking.
@@ 5.3.3 Limitations
Radiant supports domain experts with specifying patterns in sensor data that can be used to detect the execution of process activities. We do not provide any guidance on how to setup the sensor-based activity detection system including which sensors to use or how to combine them to derive process-level events. These are the responsibilities of the IoT engineer and domain expert; our goal is to enable the domain expert to efficiently formalize the sensor patterns for automated processing. The sensor patterns and conditions that are currently supported by Radiant are sufficient to cover activity detections in our laboratory setups (cf. Sections 2.1 and 2.2). We acknowledge that these are simplified setups to simulate real-word IoT systems and leave evaluations in larger scale, more realistic settings to future work. These might for example reveal limitations of the usability of the DSL and performance of the CEP engine when working with complex, high-velocity IoT sensor data, which would require additional abstractions, hierarchical patterns, and pre-processing before patterns can be specified using Radiant.
@@
A specific limitation of the pattern-based approach is that we currently analyze activity executions in isolation, i.e., we do not support multiple concurrent activities to be executed by or across the same stations involving the same sensors at a time, which may lead to a superimposition of sensor data. Furthermore, we experienced some potential issues regarding the conjunction of multiple change conditions in one pattern, which–depending on the sampling frequencies of the sensor data–may also be represented as two subsequent patterns when the changes between two subsequent events from the sensors considered in the change patterns do not happen at the same time. Finally, we did not consider more advanced patterns from the domain of CEP (e.g., time or count-based aggregations [59]) or relationships among the individual patterns (cf. Allen’s Interval Algebra [77]), yet. These extensions will be subject to future work.
@@
In general, our proposed approach is not intended to replace process mining to analyze process and activity executions, but rather to enable it by bringing the IoT sensor data to an appropriate abstraction level where subsequent process mining can provide more reasonable insights. Process mining could already be applied to analyze the low-level IoT sensor data from the sources (e.g., to discover processes). However, due to the very fine-grained measurements, the complexity of the resulting processes models and process representations is very likely to be too high for analysts to understand and efficiently analyze [19]. Finding the correct abstractions and granularity levels in this context [17] is the responsibility of the domain expert using Radiant. Note that while we are emitting process level events (in XES format [62]) to enable traditional event-centric process mining, our approach can also be adapted to object-centric event logs [78] following the models proposed in [79] and [80], which support IoT sensors as data source to be abstracted to process events and correlated with objects.
@@ 6 Conclusion and Future Work
In this work we presented the domain-specific language Radiant, implemented based on the Langium framework, for detecting process activity executions from IoT sensor data. Radiant enables domain experts to specify patterns in sensor data that indicate start, end, and intermediate points within the execution of an automated or manual activity. The Radiant applications are translated to CEP apps and deployed to the corresponding runtime system according to your proposed software architecture, which enables online pattern detection from the sensor streams and subsequent streaming process mining. In contrast to machine learning-based approaches, Radiant applications are light-weight, understandable, and capable of online activity detection from sensor streams. Domain experts created several Radiant applications to monitor activity executions in smart manufacturing and smart healthcare processes using a variety and number of sensors. The evaluation results show that a high quality of activity detection can be achieved through the domain expertise by anticipating variations in sensor data. The results also revealed improvement potential by integrating more advanced concepts into the DSL, considering the process context, and adding sensors to the IoT system.
In future work, we will move the experiments and analysis of patterns to be integrated into Radiant to more complex, real world scenarios and IoT settings. This will also include a user study with domain experts to evaluate the usability and efficiency of Radiant. Furthermore, we will use the results of the activity detections in subsequent online process mining analyses to check for process conformance at runtime and resolve potential ambiguities based on the process context. This way we will be able to provide online feedback and suggestions about process conformance to the end-users. We will also investigate the combination of ML-based approaches together with pattern specification in Radiant resulting in a hybrid activity detection system. We envision that having a robust way of detecting and representing a process activity based on sensor data will facilitate the development of digital twins of business processes enabled by the IoT [81].
@@ Acknowledgments
This work has received funding from the Swiss National Science Foundation under Grant No. IZSTZ0_208497 (ProAmbitIon project). This work has been supported by the Internet of Processes and Things (IoPT) community [^4] https://zenodo.org/communities/iopt/about through discussions, datasets, and feedback.

### v2

@@ 5.6.1 General Findings
The results of the evaluation show the validity of the proof-of-concept implementation of Radiant to be used for composing event abstraction applications, translated to corresponding CEP applications, that are capable of detecting process activity executions from IoT data. CEP proves to be a suitable technology allowing effective and efficient pattern detection in IoT data streams. However, the variations of the F1 score demonstrate an important issue we observed when working with process executions and sensors in the physical world: variations in activity executions and the associated IoT data, which show the limitations of pattern-based approaches. The knowledge of domain experts using the DSL hereby plays an essential role to influence and improve the quality of detections by anticipating these potential variations in the sensor data that have to be encoded in the Radiant patterns. For some types of activities, the results show a drop in the detection quality, which is usually due to sensor data not being sufficiently available to distinguish similar activities from each other or to detect them at all. Sections 5.6.2 and 5.6.3 discuss these aspects for the two investigated scenarios in more detail. In general, we face a trade-off between increasing the privacy-invasiveness of the monitoring setup using more capable sensors (e.g., cameras) to increase the quality of activity detections, but then also requiring more complex pre-processing steps (e.g., using computer vision) for abstracting the sensor data to be used in the pattern specifications and sacrificing runtime capabilities.
In a preliminary informal user study we asked the domain experts who created the Radiant applications for the experiments about their experience. Anecdotal evidence from the domain experts highlights the integrated support through the IDE for efficient “programming” of the activity detection patterns. The pattern specification was seen as accessible, also by non-programmers such as the involved domain experts themselves, and the sensor patterns with the link to process activities were perceived as understandable. Note that a more comprehensive user study with domain experts is required to evaluate Radiant’s suitability and usability for specifying process-level events based on IoT patterns. This is subject to future work.
@@
The seamless workflow from writing a Radiant application to the live system detecting activities from IoT sensor data allows us to conduct efficient evaluations of Radiant applications. Given recorded IoT data and a ground truth log, we are able to quickly inform the domain expert about the detection qualities of the composed applications, potential ambiguities and the need for improvements, e.g., by integrating additional patterns and installing additional sensors to increase the monitoring quality, if possible, or to constrain the variability associated with the execution of an activity. As shown with the experiments, the availability of a pre-existing, recorded IoT sensor log and associated ground truth of activity executions can also help the domain expert in a reverse engineering approach to identify relevant patterns in the sensor data that should be part of the Radiant applications. Based on these preliminary insights, we can hypothesize that Radiant, a DSL serving as an abstraction on top of event processing languages [boubeta-puig_model-driven_2014], makes event abstraction using IoT data more accessible to IoT domain experts than other pattern or ML-based approaches, because it integrates important concepts from IoT and BPM as first class citizens (cf. RQ1 and Objective 1).
Compared to ML-based models and inference, the proposed pattern-based approach is light-weight as it does not require an existing corpus of data and extensive model training (Objective 2). Nevertheless, ML-based sensor processing components can be integrated into the Radiant applications assuming that they emit pre-processed, abstracted sensor events to the message broker according to the proposed runtime architecture (cf. Section 4.6). The applications written by the domain experts and used for the experiments comprise 107 lines of Radiant code for detecting 12 different types of activities in the smart factory processes, and 224 lines of Radiant code for detecting 13 different types of activities the smart healthcare processes. The Radiant application for the smart factory generated 12 Siddhi apps, one for each activity, which are each between 35 to 45 lines of code in size, totaling to 455 lines of Siddhi code. The Radiant code for detecting the 13 activities in the smart healthcare setup translated to 13 Siddhi apps, each between 35 and 50 lines of code, totaling to 545 lines. Based on these numbers, we can observe that Radiant introduces an abstraction/reduction of the required code base by factors 2.4–4.25. On average, the file size of one generated Siddhi application to be executed on the CEP platform is 5.3 KB, indicating a relatively small disk memory footprint (Objective 2). Besides the relevant pattern in the sensor data, this also includes all the technical configurations to connect to the event sources and sinks (e.g., the MQTT message broker). As indicated in Section 3.1, a Siddhi file generated from Radiant is smaller than TinyML-models fine-tuned for activity detection in IoT deployments.
In contrast to related approaches only working in offline analysis settings, the underlying CEP engine (part of the WSO2 stream processor [fremantle2015reference]) is specifically designed to process large streams of event data and detect event patterns at runtime, allowing us to provide online feedback regarding the detected activities. The results of our experiments, the small memory and data footprints, and existing benchmarks [perera2014solving], indicate that the proposed software architecture (cf. Section 4.6) may serve as an extensible and scalable solution to process IoT data from arbitrary sensors (Objective 4) and to emit process-level events for detected activity executions beyond small scale IoT deployments (cf. RQ2 and Objective 3). By following best practices from software engineering and architecture, e.g., defining the granularity of one CEP app to match size of one activity, we achieve high cohesion and low coupling in the system, which facilitate extensibility and fault-tolerance [martin2017clean]. Note that we do not expect the activity detections to be used for real-time control of CPS or their components, but to analyze process executions and provide feedback at runtime to users (e.g., a warning in case the healthcare worker forgot to sanitize the hands before performing a specific treatment activity, indicating non-conformance with the process model [franceschetti2023proambition]) or to other information systems for further analysis or storage.
5.6.2 Findings: Smart Manufacturing
@@
The results of the activity detections in the smart manufacturing scenario (cf. Figure 9) show that we can achieve relatively high levels of correct detections when there is only one specific type of activity executed on a production station (e.g., in the sorting machine, oven, and milling machine). Here we are also able to handle variations in sensor patterns quite well via underspecification of patterns and conditions, and alternative patterns (e.g., in the sorting machine). The vacuum gripper robot (VGR) is a central transport entity, which exhibits similar sensor patterns in its movements for different types of activities. Activities can be mostly distinguished based on the x,y,z-target coordinates of the robot, which need to be discretized by the expert first to represent specific locations. A notable station is the high-bay warehouse (HBW) where there are two different types of activities executed (store and unload). The relevant movement patterns are mostly identical which leads to rather low quality detections. Here we miss a sensor to determine if the buckets being stored or unloaded contain a workpiece or not, which discerns both activities. Moreover, these two activities generally exhibit high amounts of variability as they are executed differently based on the locations of the buckets in the 3x3 storage matrix. Adding more conditions, intermediate and disjunctive patterns might help to increase accuracy here, but will also lead to complex Radiant applications and overfitting, which might decrease the detection quality in the presence of sensor variations again [seiger5165943online]. Alternatively, the granularity of the activities can be adjusted to, e.g., consider the storage and unloading of an item in each slot of the warehouse as a distinct activity type (e.g., store/unload in/from bucket 1).
@@
Note that in general, the Event Analysis metrics are better for all detections than the Two Set metrics, which also consider the specific points in time when the activity executions happened [kurz2024activity]. This misalignment can be attributed to the software stack controlling the individual production stations [seiger2022integrating]. Here we use a layered service-based architecture which introduces some latencies between service executions, logged by the BPM system, and the actual executions in the IoT system that manifest themselves as changes in its sensors and actuators. These latencies might be unintentional (e.g., due to network communication) or intentional (e.g., the services might perform some calculations or data processing before starting to manipulate the sensors and actuators).
5.6.3 Findings: Smart Healthcare
@@
The results of the activity detection in the healthcare scenario (cf. Figure 10) show a lower quality of detections. As described in Section 2.4, the activities in this scenario are almost completely manual and strongly affected by variations in how they are executed, and in the underlying IoT sensor data, which cannot be completely handled by discretizations and alternative detection patterns. While we achieve very good detections for the Sanitize hands activity relying on sensor patterns in the scale and distance sensor in front of it (cf. Figure 8), activities related to the interaction with the donor (e.g., apply/remove tourniquet, insert needle, monitor patient) are harder to detect and distinguish from each other as their executions are quite similar and rely on the same sensors and similar patterns (cf. simultaneous detection of low-level patterns for activities in Figure 8). For these activities, we currently consider using cameras in combination with the existing sensors for disambiguation [franceschetti2025proambition]. Moreover, the starting and stopping of the blood drawing machine rely on the same button to be pressed, thus we cannot distinguish them from each other purely on the sensor data. Currently, we also investigate considering the process context (e.g., which activities happened before) for further disambiguation. Moreover, similar to the event log serving as ground truth for the smart manufacturing settings, we also observe timeshifts when aligning the start and end of activities for the Two Set metrics in the smart healthcare setting, which leads to a decreased performance compared to Event Analysis. These shifts can be attributed to the creation of the ground truth event log, which is based on manual monitoring and tracking by an external observer of the executed processes. Here, the detection of start and end patterns from the IoT sensor data might not always be aligned with the manual tracking.
5.6.4 Comparison with Related Approaches
We are aware of three related approaches that follow the same functional objective in proposing event abstraction mechanisms to detect activity executions in similar IoT settings and based on similar IoT data. In [fi15020077], the authors propose a method for domain experts to identify activity executions including the relevant change patterns in the IoT data, and to then manually annotate the data with process level events (i.e., start and end of an activity). This approach relies entirely on the domain expert to manually analyze a given dataset for activity executions (Objective 1), which is not feasible in settings with a high number of sensors and activities that have to be identified and labeled in an unknown dataset.
In [seiger2025case], the authors propose to use temporal case-based reasoning for the classification of unknown IoT data according to a known set of activity–IoT data correlations (cases). This approach requires the domain experts to specify similarity measures for the sensors to be used for the matching with the cases. The authors demonstrate that activity detections in the smart factory scenario are possible and feasible (average F1 score of 0.837). While these numbers are an improvement over our experiments, the approach assumes that the IoT data to be classified has already been pre-segmented according to its start and end, and it only works in offline analysis (Objective 2). The underlying models in the case base are more complex than the Radiant applications, but still comprehensible (Objective 1) and relatively light-weight (Objective 3). The approach is applicable to arbitrary sensor data (Objective 4).
In [garcia2025semi], the authors propose a sensor data analysis pipeline consisting of clustering and sequence mapping to transform symbolic sequences representing the activity executions into automata, which are able detect the occurrences of the encoded sequences. In the last step, the authors rely on the domain expert to confirm or manually adjust the identified clusters of sensor data and to provide a label that associates the data with an executed activity. Thus, the domain expert is still involved and high F1 scores can be achieved via the manual adjustments. However, the understandability is decreased (Objective 1) due to the clustering and encoding of the multi-variate timeseries of sensor values. The approach is more data and processing intensive (Objective 2) and not capable of online event abstraction (Objective 3).
@@ 5.6.5 Limitations
Radiant supports domain experts with specifying patterns in sensor data that can be used to detect the execution of process activities. We do not provide any guidance on how to setup the sensor-based activity detection system including which sensors to use or how to combine them to derive process-level events. These are the responsibilities of the IoT engineer and domain expert; our goal is to enable the domain expert to formalize the sensor patterns for automated processing. The sensor patterns and conditions that are currently supported by Radiant are sufficient to cover activity detections in our laboratory setups (cf. Sections 2.3 and 2.4) as discussed with the involved domain experts. We acknowledge that these environments are simplified setups to simulate real-word IoT systems and leave evaluations in larger scale, more realistic settings to future work. These might for example reveal limitations of the usability of the DSL and performance of the CEP engine when working with complex, high-velocity IoT sensor data, which would require additional abstractions, hierarchical patterns, and pre-processing before patterns can be specified using Radiant. As a specific threat to validity we emphasize again that a formal validation of Radiant’s usability has not been conducted so far. Only through this study, which is part of future work, we can evaluate that Radiant does not only support domain experts with specifying sensors patterns for process event abstraction, but that it also does so in an efficient, learnable and usable manner.
@@
Another specific limitation of the pattern-based approach is that we currently analyze activity executions in isolation, i.e., we do not support multiple concurrent activities to be executed by or across the same stations involving the same sensors at a time, which may lead to a superimposition of sensor data. Furthermore, we experienced some potential issues regarding the conjunction of multiple change conditions in one pattern, which–depending on the sampling frequencies of the sensor data–may also be represented as two subsequent patterns when the changes between two subsequent events from the sensors considered in the change patterns do not happen at the same time. Finally, we did not consider more advanced patterns from the domain of CEP (e.g., stateful patterns, time or count-based aggregations [etzion2010event]), or relationships among the individual patterns (cf. Allen’s Interval Algebra [allen1983maintaining, barricelli_visual_2017]) because they did not emerge as requirements in the presented IoT scenarios and also may refer to subsequent process mining analyses, rather than event abstraction.
@@
In general, our proposed approach is not intended to replace process mining to analyze process and activity executions, but rather to enable it by bringing the IoT sensor data to an appropriate abstraction level where subsequent process mining can provide more reasonable insights. Process mining could already be applied to analyze the low-level IoT sensor data from the sources (e.g., to discover processes). However, due to the very fine-grained IoT-based measurements, the complexity of the resulting processes models and process representations is very likely to be too high for analysts to understand and efficiently analyze [brzychczy2025process]. Finding the correct abstractions and granularity levels in this context [zerbato2021granularity] is the responsibility of the domain expert using Radiant. Note that while we are emitting process level events (in XES format [gunther2014xes]) to enable traditional event-centric process mining, our approach can also be adapted to object-centric event logs [van2019object] following the models proposed in [bertrand2023nice] and [bertrand2025objectcentriccoremetamodeliotenhanced], which support IoT sensors as data source to be abstracted to process events and correlated with objects.
@@ 6 Conclusion and Future Work
In this work we presented the domain-specific language Radiant, implemented based on the Langium framework, for detecting process activity executions from IoT sensor data. Radiant allows domain experts to specify event abstractions patterns in sensor data that indicate start, end, and intermediate points within the execution of an automated or manual activity. The Radiant applications are translated to CEP applications and deployed to the corresponding runtime system according to your proposed software architecture, which enables online pattern detection from the sensor streams and subsequent streaming process mining. In contrast to machine learning-based approaches, Radiant applications are light-weight by focusing on sensor patterns not requiring existing data, understandable and easily modifiable by making the sensor patterns explicit, and capable of online activity detection from sensor streams enabled by the software architecture. Domain experts without programming skills created several Radiant applications to monitor activity executions in smart manufacturing and smart healthcare processes using a variety and number of sensors. The evaluation results show that the CEP apps generated from the Radiant applications can be effectively used to detect process activity executions at scale. Variations in the sensor data and activity execution have a strong influence on the detection quality, which needs to be anticipated by domain expert in the pattern definitions. The results also revealed improvement potential by integrating more advanced concepts into the DSL, considering the process context, and adding sensors to the IoT system.
In future work, we will move the experiments and analysis of patterns to be integrated into Radiant to more complex, real world scenarios and IoT settings. This will also include a more comprehensive user study with domain experts to evaluate the usability and efficiency of Radiant, which has not been evaluated in terms of usability so far. Furthermore, we will use the results of the activity detections in subsequent online process mining analyses to check for process conformance at runtime and resolve potential ambiguities based on the process context. This way we will be able to provide online feedback and suggestions about process conformance to the end-users. We will also investigate the combination of ML-based approaches together with pattern specification in Radiant resulting in a hybrid activity detection system. We envision that having a robust way of detecting and representing a process activity based on sensor data will facilitate the development of digital twins of business processes enabled by the IoT [fornari2024digital].
@@ Acknowledgments
This work has received funding from the Swiss National Science Foundation under Grant No. IZSTZ0_208497 (ProAmbitIon project). This work has been supported by the Internet of Processes and Things (IoPT) community through discussions, datasets, and feedback.